OpenStudy (anonymous):

a right triangle has one leg measuring 8 in and the hypotenuse measuring 12 in. find the length of the other leg. express your answer in simplest radical form.

OpenStudy (anonymous):

It's about 8.9

OpenStudy (ikram002p):

 apply Pythagorean Theorem 12^2=8^2+x^2 x^2=80

OpenStudy (anonymous):

8.9 equals x

OpenStudy (ikram002p):

x^2-80 =0 use the general formula of quadratic equation

OpenStudy (ikram002p):

@Nathan28 ur answer is correct but the qs want the answer as simplest radical form. which is a/b form not sqrt 80 or decimal fraction